Use the substitution theorem evaluate,


Integral 0 to 2 t^2[1+t^3]^-1/2dt

1
Expert's answer
2022-03-29T11:46:13-0400

"I=\int_0^2 \frac{t^2}{\sqrt{1+t^3}}dt"

Substitution:

"u=t^3+1" ; "du=3t^2dt";

"u_1=0^3+1=1" ; "u_2=2^3+1=9".

"I=\int_0^2 \frac{t^2}{\sqrt{1+t^3}}dt=\frac13\int_0^2 \frac{3t^2dt}{\sqrt{1+t^3}}=\frac13\int_1^9 \frac{du}{\sqrt{u}}=""\frac23\sqrt u|_1^9=\frac23(3-1)=\frac43"